SPECTROSCOPIC CHARACTERIZATION OF HIGHLY PREDISSOCIATIVE LEVELS OF THE OH A^{2}Σ^{+} STATE

\underline{LOGAN P. DEMPSEY}, ERIKA L. DERRO, ILANA B. POLLACK, and MARSHA I. LESTER, Department of Chemistry, University of Pennsylvania, Philadelphia, PA 19104-6323.

The OH radical plays a major role in atmospheric and combustion environments, where it is generally detected by Laser Induced Fluorescence (LIF) using the A^{2}Σ^{+} − X^{2}Π band system. Excited vibrational levels of the OH A^{2}Σ^{+} (v ≥ 3) state have been difficult to study by LIF due to rapid predissociation, which results in a fluorescence quantum yield that is essentially zero. In the present work, the OH A^{2}Σ^{+} (v = 4) state is characterized using a Fluorescence Depletion Infrared (FDIR) spectroscopy technique. A UV laser promotes OH radicals from the v′′ = 1 level of the ground X^{2}Π state to the v′ = 2 level of the excited A^{2}Σ^{+} state. An IR laser further excites the OH radicals on an overtone transition to the v = 4 level of the A^{2}Σ^{+} state. Fluorescence is collected from OH A^{2}Σ^{+} (v′ = 2) exclusively, which is depleted when the IR laser is resonant with a 4 ← 2 overtone transition. The IR frequencies of various P- and R-branch lines of the overtone transition are measured, and spectroscopic constants for the A^{2}Σ^{+} (v = 4) state are extracted using combination differences. The rotational, centrifugal distortion, and spin-rotation constants are determined, along with the vibrational frequency for the overtone transition. These constants are compared with calculated values from Dunham coefficients for the OH A^{2}Σ^{+} state and a merged fit of experimental data from previous studies.